Crew 239 Sol Summary for November 8th

Sol: 1

Summary Title: First Day of Martian Activity!

Author’s name: Madison Rooth

Mission Status: Nominal

Sol Activity Summary: Our crew awoke this morning and shared a breakfast of delicious pancakes cooked by Crew Artist Brian Murphy. After this, EVA Team 1 of CO Keith Crisman, XO Madison Rooth, and MCA Brian Murphy embarked to Marble Ritual for a geologic sampling and training EVA at 1000.Team 1 arrived back at the Hab at 1059 and cooled off while preparing EVA Team 2 for their mission to Marble Ritual. EVA Team 2 left Hab at 1332 and consisted of GHO Cynthia Montanez, HSO Emily Milne, and ENG Aidan Moga. Following EVA Team 2’s return to the Hab at 1500, we shared a delicious meal of rice, vegetables, and chicken. ENG Aidan and CO Crisman attempted to fix our waste facility (toilet) due to a freshwater leak). The crew then worked on individual research until dinnertime.

Look Ahead Plan: Conduct EVAs to Kissing Camel Ridge

Anomalies in work: No anomalies

Weather: Sunny and hot

Crew Physical Status: All good

EVA: 2 to Marble Ritual

Reports to be filed: Operations report on it’s way!

Support Requested: None at this time

Crew 239 Sol Summary for November 7th

Crew 239 Sol Summary Report 11-07-2021

Sol: 1

Summary Title: Beginning of sim

Author’s name: Madison Rooth

Mission Status: Nominal

Sol Activity Summary: We arrived in Hollow Mountain to fuel the CrewCar at 10:55am and arrived at the Hab at 11:45am. We conducted an overview of the site and met with Shannon shortly after. After system introductions and acclimating the Crew to their new surroundings, two members, XO Madison Rooth and ENG Aidan Moga followed Shannon into Hanksville to fill the HabCar tank with water from Hollow Mountain. Due to difficulties with the hose and pump, it took several hours to pump the water into the Static Tank. After this was completed, HSO Emily Milne and GHO Cynthia Montanez made the second trip into town to fill the HabCar tank and transfer it to the Static Tank. CO Keith Crisman, MCA Brian Murphy, and XO Madison Rooth repaired the window in the Hab’s main door due to it breaking from the high winds.

Look Ahead Plan: Conduct Marble Ritual EVAs

Anomalies in work: No anomalies

Weather: Sunny and windy

Crew Physical Status: All good

EVA: No EVA

Reports to be filed: HSO Report

Support Requested: None at this time

Crew 260 Sol Summary for November 4th

Crew 260 Sol Summary Report 04-11-2021

Sol: 3

Summary Title: The end of sim

Author’s name: Jen Carver-Hunter

Mission Status: Nominal

Sol Activity Summary: After eating our last breakfast in sim, we returned to Earth for a final lesson with Dr Rupert and Atila about biosignatures. Following our lesson, we went out into the “field” to the east of the Hab to look for biosignatures. We specifically focused on finding desert varnish and lichen. After returning to the Hab, we took a group field trip into town to visit Kathy at the rock shop, and enjoyed burgers and shakes at Stan’s. Returning to the Hab for our last afternoon together was bittersweet. We donned flight suits one last time to launch rockets. Then several members of the crew returned to road 1101 to search for the elusive purple galaxy rock. Upon their return, we began our end of mission duties – cleaning and writing reports.

Look Ahead Plan: Finish cleaning and check out

Anomalies in work: No anomalies

Weather: Sunny, warm, and pleasant

Crew Physical Status: All good

EVA: No official EVA

Reports to be filed: Operations report, Journalist report, end of mission summary

Support Requested: none at this time

Crew 228 Sol 11 Summary October 8th

Sol: 11

Summary Title: The Mission Winds Down

Author’s name: Dave Laude

Mission Status: Nominal

Sol Activity Summary:

After breakfast discussed the day’s activities, mostly cleaning and preparing campus for our departure.

Lindsay continued metagenomic regolith experiments in the ScienceDome and
Inga prepared for her upcoming presentation at the Mars Society
Conference. She will present machine learning findings about social
dynamics in space analogs.

For a Show and Tell during lunch break Dave talked about radio in the early 1920s and demonstrated a 1924 battery operated 2-tube receiver.

Dave and Inga took food inventory and entire crew performed general cleaning.

Weather: Cloudy, windy, dark, light rain and cold. Nice day to stay inside, clean and write reports.

Crew Physical Status: All crew nominal, except one scraped back of hand on irregular piece of wood. Very little slow bleeding stopped itself in a few minutes. Wound cleaned and bandaged.

EVA: None

Reports to be filed: Commander, Science, Photo, Sol Summary, Journalist.

Support Requested: None

END

Crew 228 Sol Summary October 6th

Sol: 9

Summary Title: Sim Nears End

Author’s name: David Laude (XO)

Mission Status: Nominal, Hab fully prepared for in-sim operations

Sol Activity Summary: The day started with breakfast pancakes made
by Dave, during which time was discussed how our daily routine will run for
the next few sols. Dave and Inga completed EVA #6 to Candor Chasma. Jin started working
on a special edition of the journalist report for the final soul. Jin and Inga also
have been working on chili for dinner. Some greens were harvested from the
GreenHab for consumption. Lindsay has been using a DNA sequencer to identify
organisms. She investigated two regolith samples and got very low DNA yield so far.
She will continue with additional samples to see if she can get better DNA yield.
Either way, she will determine if she can identify particular microbes from low DNA yield.

Look Ahead Plan: Tomorrow will have a 7-8 hour break in sim for Dave and Jin as Jin needs to take a Covid test ahead of his flight from Grand Junction to Canada and Dave with a US drivers license will drive him for the test. Travel will be by Crew Car to Fruita just west of Grand Junction. The crew’s sim will soon end as they will clean the station for the next crew this Friday and depart MDRS Saturday for home.

Anomalies in work: Radio failure in field issues mostly battery related. Crew Engineer Dave will conduct more battery tests in hopes of culling out some weak batteries. Tests are done by charging a battery and then with a radio set to high power transmit level the push to talk button is held down with a clamp. A timer is used to time when the battery indicator in the radio indicates a depleted battery.

Weather: High of 71F today, partly cloudy. increasing winds in the afternoon with gusts ~20-25 MPH.

Crew Physical Status: All crew nominal.

EVA: Eva #6 with Dave and Inga to Candor Chasma was fully successful. Returned to airlock within seconds of planned arrival.

Reports to be filed: Science, Photo, Sol, Journalist

Support Requested: None

END